A monopolist sells a product in 2 neighboring states and practices 3rd degree price discrimination by charging separate prices in each state. The market demand curve in state 1 is given by: q1=50-p1 and by q2=90-1.5p2 in state 2. The firm can produce the product at MC=10 in both states. A. What price should the monopolist charge in each state to maximize total profit? B. Suppose that a reseller could purchase the product in the low price state and transport it to the high price state at a cost of $4 per unit. How would the reseller influence the monopolist's pricing strategy in the high price state?
1
Expert's answer
2019-07-30T09:25:19-0400

A. Profit is maximized when MR = MC.

Q1 = 50 - p1 or p1 = 50 - Q1, so MR1 = TR1' = 50 - 2Q1.

50 - 2Q1 = 10,

Q1 = 20 units,

p1 = 50 - 20 = 30.

Q2 = 90 - 1.5p2 or p2 = 60 - Q2/1.5, so MR2 = TR2' = 60 - 4/3Q2.

60 - 4/3Q2 = 10,

Q2 = 37.5 units,

p2 = 60 - 37.5/1.5 = 35.

B. If a reseller could purchase the product in the low price state and transport it to the high price state at a cost of $4 per unit, then the reseller would decrease the monopolist's price in the high price state.